OpenStudy (anonymous):

Find 3 1/2 divided by 2 3/4. Write in simplest form.

OpenStudy (whpalmer4):

First, write each mixed number as a fraction. Second, recall that dividing fractions can be accomplished by multiplying by the reciprocal of the divisor, which is just flipping the divisor over. For example, \[\frac{3}{2} / \frac{1}{2} = \frac{3}{2}*\frac{2}{1} = 3\]

OpenStudy (anonymous):

I'm confused.

OpenStudy (whpalmer4):

Okay, 1 step at a time. Do you know how to convert 3 1/2 to a fraction?

OpenStudy (anonymous):

Do you do 3 x 3 + 1?

OpenStudy (whpalmer4):

Not quite. \[3 \frac{1}{2} = 3*\frac{2}{2} + \frac{1}{2} = \frac{6}{2} + \frac{1}{2} = ?\]

OpenStudy (anonymous):

The problem is 3 1/3 divided by 2 3/4 how did hyou get 3 1/2

OpenStudy (whpalmer4):

\[\frac{6}{2} + \frac{1}{2} = \frac{7}{2}\] What you maybe meant to do was \[3\frac{1}{2} = \frac{3*2+1}{2} = \frac{7}{2}\]

OpenStudy (whpalmer4):

Go look at what you wrote up top!

OpenStudy (anonymous):

Oops I meant 3 1/3

OpenStudy (whpalmer4):

Okay, so then yes, that would be 10/3 for the first fraction. What about the 2 3/4, what is that as a fraction?

OpenStudy (anonymous):

It would be 11/4

OpenStudy (whpalmer4):

Right. So our problem is \[\frac{10}{3}/ \frac{11}{4}\] As I mentioned earlier, you can divide two fractions by inverting the second fraction and multiplying. \[\frac{10}{3} / \frac{11}{4} = \frac{10}{3} * \frac{4}{11} = ?\]

OpenStudy (anonymous):

40/33?

OpenStudy (whpalmer4):

Correctamundo! :-)

OpenStudy (anonymous):

Yay! I need to make it not improper I think.

OpenStudy (whpalmer4):

You could do that. What do you get?

OpenStudy (anonymous):

1 6/33

OpenStudy (whpalmer4):

Close, but not quite. 40 - 33 = 6?

OpenStudy (anonymous):

What is it? I'm so confused lol.

OpenStudy (whpalmer4):

\[\frac{40}{33} = \frac{33+7}{33} = \frac{33}{33} + \frac{7}{33} = 1\frac{7}{33}\]

OpenStudy (anonymous):

ohh ok!

OpenStudy (anonymous):

Find 3 5/6 - 5 2/3. Write in simplest form.

OpenStudy (whpalmer4):

Do you know how to find common denominators? You'll want to convert those two mixed numbers to fractions, find the common denominator, then subtract.